# Beyond Sharp Minima: Robust LLM Unlearning via Feedback-Guided Multi-Point Optimization

### v1

@@ Abstract
Current LLM unlearning methods face a critical security vulnerability that undermines their fundamental purpose: while they appear to successfully remove sensitive or harmful knowledge, this “forgotten” information remains precariously recoverable through relearning attacks. We identify that the root cause is that conventional methods optimizing the forgetting loss at individual data points will drive model parameters toward sharp minima in the loss landscape. In these unstable regions, even minimal parameter perturbations can drastically alter the model’s behaviors. Consequently, relearning attacks exploit this vulnerability by using just a few fine-tuning samples to navigate the steep gradients surrounding these unstable regions, thereby rapidly recovering knowledge that was supposedly erased. This exposes a critical robustness gap between apparent unlearning and actual knowledge removal. To address this issue, we propose StableUN, a bi-level feedback-guided optimization framework that explicitly seeks more stable parameter regions via neighborhood-aware optimization. It integrates forgetting feedback, which uses adversarial perturbations to probe parameter neighborhoods, with remembering feedback to preserve model utility, aligning the two objectives through gradient projection. Experiments on WMDP and MUSE benchmarks demonstrate that our method is significantly more robust against both relearning and jailbreaking attacks while maintaining competitive utility performance.
@@ 2.1 Machine Unlearning in Large Language Models
Against this backdrop, this paper focuses on how to improve a representative set of approximate weight-based unlearning techniques that have become the baselines in recent LLM studies, including: i). Gradient Ascent on $D_{f}$ (GA) (Jang et al., 2022), ii). Gradient Ascent on $D_{f}$ + Gradient Descent on $D_{r}$ (GA + GD) (Liu et al., 2025; Maini et al., 2024), iii). Gradient Ascent on $D_{f}$ + KL Divergence on $D_{r}$ (GA + KL) (Maini et al., 2024; Shi et al., 2024), iv). Negative Preference Optimization on $D_{f}$ (NPO) (Zhang et al., 2024) and v). Representation Misdirection for Unlearning (RMU) (Li et al., 2024). They are detailed in Appendix A.2.
@@
Recent audits reveal that LLMs subjected to “unlearning” procedures still leak private data. Membership-inference accuracy often rebounds on “forgotten” samples (Duan et al., 2024), and prompt-injection (Shumailov et al., 2024) can smuggle hidden instructions past guardrails, reviving passage-level content the model was meant to forget. Even benign fine-tuning on a handful of public texts can quickly restore excised knowledge (Hu et al., 2024). To counter this, Xu et al. (2025b) combines masking, distillation, and fact regularization to resist inference; min–max objectives enhance robustness (Fan et al., 2025); and random noise-based augmentation reframes forget tokens as back-door triggers to neutralize them (Huu-Tien et al., 2025). Privacy-oriented work leverages token-specific training to suppress memorized tokens, halving membership-inference success rates with minimal utility cost (Tran et al., 2025), and fairness-centric studies show that reducing variance gaps likewise mitigates adversarial recovery of forgotten data (Tran & Woo, 2025).
@@ 3.1 LLM Unlearning
We begin by formulating LLM unlearning. Let $D_{t}$ denote the original training dataset, from which a pretrained original model $f_{\text{O}}(\cdot;\theta)$ is obtained, where $\theta$ is the parameters of $f_{\text{O}}$. The goal of an unlearning algorithm $\mathcal{MU}$ is to remove the influence of a designated forget set $D_{f}$ from $f_{\text{O}}$ (Liu et al., 2025; Geng et al., 2025; Liu et al., 2024b). To ensure that the unlearning process does not significantly degrade the model’s overall utility, a retain set $D_{r}$ is typically introduced (Ren et al., 2025; Ji et al., 2024). In practice, the retain set $D_{r}$ and the forget set $D_{f}$ are disjoint, i.e., $D_{r}\cap D_{f}=\emptyset$. Based on $D_{f}$ and $D_{r}$, $\mathcal{MU}$ typically defines two loss terms: a forget loss that penalizes residual influence from $D_{f}$, and a retain loss that encourages preservation of performance on $D_{r}$. These objectives capture the dual goals of forgetting and retention, and can be expressed as the following regularized optimization problem (Pan et al., 2025):
@@ 3.2 Relearning Attacks in LLM Unlearning
The robustness issue of LLMs in the context of unlearning is primarily reflected in the vulnerability of current machine unlearning methods to relearning attacks (Fan et al., 2025). These attacks aim to rapidly recover deleted knowledge by performing lightweight fine-tuning on the unlearned model $f_{\text{U}}$ using only a small number of samples from the forget set. The attacker’s objective is as follows:
@@
where $\delta$ denotes the adversarial update to the $f_{\text{U}}$’s parameters, $D_{sub_{f}}\subset D_{f}$ is a small subset of the forget set used in the attack, $\ell_{\text{relearn}}$ is the relearning objective, which is often defined to counteract the unlearning process, such as the general fine-tuning loss or the negative of the forget loss.
@@ 4.1 Motivation
To illustrate the robustness vulnerability in current LLM unlearning methods and establish the foundation for our approach, we conduct a systematic analysis of the relearning attack. We first perform unlearning on the Zephyr-7B-beta model (Tunstall et al., 2023) using the standard Gradient Ascent (GA) method on the WMDP Bio dataset (Li et al., 2024). After unlearning, we simulate an adversary who fine-tunes the unlearned model for only two epochs on small subsets of the original forget set $D_{f}$ (20 samples). Performance is evaluated on the WMDP Bio QA test set, where lower accuracy indicates stronger unlearning. As shown in Figure 1(a), GA-based unlearning reduces accuracy from 64.45% (original model) to 27.10%, suggesting effective suppression of target knowledge. However, the relearning
@@
When $S_{\delta}(\theta)$ is large, even tiny parameter updates can drastically change the loss. In such sharp regions, adversaries need only a few fine-tuning steps on small subsets of $D_{f}$ to reverse unlearning, directly explaining the effectiveness of small-shot relearning attacks. We visualize the loss surface around unlearned parameters by scanning two orthogonal directions $\mathbf{r}_{1},\mathbf{r}_{2}$ and plotting $z=\ell(\theta+\alpha\mathbf{r}_{1}+\beta\mathbf{r}_{2})$. As shown in Figure 2, GA-based unlearning forms a sharp basin, while our method produces a flatter basin with lower neighborhood sensitivity, thereby resisting small-shot relearning. This analysis highlights the need to explicitly control $S_{\delta}(\theta)$ during unlearning:
@@ 4.2 Overview
To address the robustness limitations of current LLM unlearning methods while explicitly balancing the dual objectives of forgetting and remembering, we propose a feedback‑guided unlearning framework. By injecting task‑specific guidance signals into the optimization loop of mainstream LLM unlearning algorithms, the overall training procedure becomes more robust and stable. Specifically, we design two kinds of feedback as follows:
  • The first is forgetting feedback. We introduce a robustness‑oriented feedback signal that exposes the model to a family of parameter perturbations. Motivated by prior work showing that relearning attacks essentially act as weight‑space disturbances (Fan et al., 2025), we require the unlearned model $f_{u}$’s performance on the forget dataset $D_{f}$ to remain invariant to small perturbations on the parameter space, thereby enhancing its resistance to relearning attacks.
@@
Inspired by meta-learning (Andrychowicz et al., 2016; Wang et al., 2020; 2021), we formulate our approach as a bi-level optimization problem: an inner-loop unlearning step produces a temporary model, while an outer-loop feedback step refines the final update direction. This formulation introduces two additional stages into the standard unlearning loop, namely the unlearning-tuning stage and the feedback stage. For the unlearning-tuning stage, we create a temporary model $f_{\text{tmp}}(\cdot;\theta^{\tau})$ by running one gradient update with the unlearning loss (e.g., GA, NPO, RMU), which provides the basic direction to forget. Our feedback signals build on top of it. For the feedback step, the temporary model is then probed by the two complementary feedback signals introduced above, generating loss terms $\mathcal{L}_{f}$ and $\mathcal{L}_{r}$, respectively. Depending on the feedback target, the two signals respectively reflect the model’s forgetting performance within its current neighborhood and utility, so it reveals how well it robustly forgets $D_{f}$ and retains knowledge from $\hat{D}_{r}$. After these two stages, the algorithm executes the final cumulative parameter update using the harmonized gradient direction.
Given that the forgetting objective $\mathcal{L}_{f}$ and remembering objective $\mathcal{L}_{r}$ can induce conflicting update directions, we further introduce a gradient harmonization strategy to resolve this dilemma. Specifically, a simple projection operation removes the component of one gradient that conflicts with the other, producing a unified update direction that simultaneously promotes thorough deletion of $D_{f}$ and faithful preservation of essential knowledge in $D_{r}$. In the following sections, we detail the complete framework and demonstrate how we improve the robustness of LLM unlearning.
@@ 4.3 Forgetting Feedback
To achieve a more thorough forgetting of the forget dataset $D_{f}$, especially to enhance the robustness of the model against relearning attacks, we propose forgetting feedback by applying various perturbation techniques to the model parameters to simulate the relearning attacks, ensuring the unlearned model’s performance on $D_{f}$ remains stable. The detailed procedure is as follows: As shown in Figure 3(a), the first step of forgetting feedback is the unlearning-tuning step. Initially, we construct a temporary model $f_{\text{tmp}}(\cdot;\theta^{\tau})$ by performing one gradient update using a standard forgetting loss (e.g., GA, GA+GD, GA+KL, NPO, RMU) on the original model $f_{\text{O}}$’s parameters $\theta$:
@@
The first category is chosen because it deliberately pushes the model toward worst-case directions, thereby simulating adversarial attempts to “re-awaken” forgotten information. This group includes Sharpness-Aware Perturbation (SAP(Foret et al., 2020), which perturbs parameters along the normalized gradient direction to capture the steepest ascent of local sharpness. It explicitly targeting the direction that maximally increases loss, which is a typical adversarial move; Gradient-Aligned Perturbation (GAP(Moosavi-Dezfooli et al., 2019), which scales the gradient to drive the model into high-curvature regions; and Historical Weight Smoothing (HWS(Izmailov et al., 2018), which averages the current weights with several past checkpoints to test whether smoothing can implicitly roll parameters back toward memorized regimes. The second category introduces stochastic disturbances, represented by Gaussian Parameter Noise (GPN(Cohen et al., 2019), which adds unbiased Gaussian noise to the parameters and serves as a non-adversarial baseline.
Together, these perturbations form a spectrum from random to adversarial, providing a comprehensive probe of how resilient the unlearned model remains to relearning attempts. Their mathematical details are provided in Appendix A.3. Then we get the forgetting feedback. In each iteration, we randomly select $T$ perturbation methods from the above candidates to generate corresponding perturbed parameters $\{\theta^{\tau^{\prime}_{i}}\}_{i=1}^{T}$. These perturbed models are then evaluated on the relevant dataset(s) according to the specific $\mathcal{MU}$ method being used. The final forgetting feedback loss is as below:
@@ 4.4 Remembering Feedback
To prevent unintended loss of useful knowledge in $D_{r}$ and preserve downstream utility, we introduce remembering feedback, which parallels the forgetting branch but focuses on retention. After the standard unlearning tuning step (Eq. 5), we obtain a temporary model $\theta^{\tau}$. We then conduct a feedback-evaluation step on a query set $Q$, randomly sampled from a small retained subset $\hat{D}_{r}\subset D_{r}$ or a public corpus, to monitor utility degradation. The cross-entropy over the $M$ queries is averaged to form the remembering feedback loss, which penalizes drops in performance and guides optimization toward preserving generalizable knowledge while erasing $D_{f}$:
@@
where $\lambda_{r}$ modulates the strength of the remembering feedback. The formulation enables our feedback-guided remembering mechanism to penalize any degradation on informative samples.
@@
At each training step, our framework produces two distinct feedback signals: a robustness-oriented forgetting loss $\mathcal{L}_{fb}$ and a utility-preserving remembering loss $\mathcal{L}_{fb_{2}}$. Combining these signals with the base forgetting objective yields the overall optimization target:
@@
We first compute their inner product $\langle g_{f},g_{r}\rangle=g_{f}^{T}g_{r}$. If this value is negative, the two directions conflict, and we project $g_{f}$ onto the sub-space orthogonal to $g_{r}$; otherwise, we keep $g_{f}$ unchanged:
@@ 5.1 Experiment Setup
We conduct experiments on two LLM unlearning benchmarks: (1) The WMDP benchmark (Li et al., 2024), which evaluates unlearning capabilities in hazardous domains such as biosecurity (WMDP-bio), cybersecurity (WMDP-cyber), and chemical safety (WMDP-chem). We primarily focus on the first two; (2) The MUSE benchmark (Shi et al., 2024), which contains two unlearning scenarios: News and Books. The former seeks to unlearn knowledge related to BBC news articles, while the latter aims to unlearn text fragments from the Harry Potter book series. Following previous literature and to demonstrate the effectiveness of our method across different models, we employ Zephyr-7B-beta (Tunstall et al., 2023) as the initial model for WMDP-bio, Mistral-7B (Jiang et al., 2023) fine-tuned on cybersecurity datasets as the initial model for WMDP-cyber, LLaMA-2 7B (Touvron et al., 2023) fine-tuned on BBC news for the News task, and ICLM 7B (Fan et al., 2024) fine-tuned on Harry Potter books for the Books task.
@@
As shown in §2.1, we use 5 methods: GA, GA+GD, GA+KL, NPO and RMU. More details are included in Appendix A.2. The performance of LLM unlearning is evaluated through $\mathcal{MU}$ Effectiveness and model Utility Retention. For WMDP, $\mathcal{MU}$ effectiveness is measured by accuracy on the WMDP test set, where lower accuracy indicates better unlearning performance. Utility retention is assessed through zero-shot accuracy on MMLU, while higher utility scores reflect better retention of general capabilities. For the MUSE dataset, following (Shi et al., 2024), we measure performance through Verbatim Memory (VerbMem) and Knowledge Memory (KnowMem) on $D_{f}$, where lower values indicate better unlearning effectiveness. The former compares the average ROUGE-L F1 score (Klimt & Yang, 2004) between model-generated continuations and ground-truth continuations for unlearning samples, while the latter compares the average ROUGE score between model responses and ground-truth answers for question-answer pairs in the $D_{f}$. Utility is computed through KnowMem on the retained set, calculated as the average ROUGE score for question-answer pairs on the retained set. We further evaluate the robustness under two adversarial scenarios: 1) Relearning attacks (Hu et al., 2024): This constitutes our primary research focus. We randomly sample relearning data from the $D_{f}$, with results averaged over 5 independent random trials. 2) Jailbreaking attacks (Ma et al., 2024; Łucki et al., 2025; Thompson & Sklar, 2024): We employ an enhanced GCG algorithm (Łucki et al., 2025) to achieve unlearning knowledge extraction through the generation of adversarial prefixes.

### v3

@@ Abstract
Current LLM unlearning methods face a critical security vulnerability that undermines their fundamental purpose: while they appear to successfully remove sensitive or harmful knowledge, this “forgotten” information remains precariously recoverable through relearning attacks. We identify that the root cause is that conventional methods optimizing the forgetting loss at individual data points will drive model parameters toward sharp minima in the loss landscape. In these unstable regions, even minimal parameter perturbations can drastically alter the model’s behavior. Consequently, relearning attacks exploit this vulnerability by using just a few fine-tuning samples to navigate the steep gradients surrounding these unstable regions, thereby rapidly recovering knowledge that was supposedly erased. This exposes a critical robustness gap between apparent unlearning and actual knowledge removal. To address this issue, we propose StableUN, a bi-level feedback-guided optimization framework that explicitly seeks more stable parameter regions via neighborhood-aware optimization. It integrates forgetting feedback, which uses adversarial perturbations to probe parameter neighborhoods, with remembering feedback to preserve model utility, aligning the two objectives through gradient projection. Experiments on WMDP and MUSE benchmarks demonstrate that our method is significantly more robust against both relearning and jailbreaking attacks while maintaining competitive utility performance.
@@ 2.1 Machine Unlearning in Large Language Models
Against this backdrop, this paper focuses on how to improve a representative set of approximate weight-based unlearning techniques that have become the baselines in recent LLM studies, including: i). Gradient Ascent on $D_{f}$ (GA) (Jang et al., 2022), ii). Gradient Ascent on $D_{f}$ + Gradient Descent on $D_{r}$ (GA + GD) (Liu et al., 2025; Maini et al., 2024), iii). Gradient Ascent on $D_{f}$ + KL Divergence on $D_{r}$ (GA + KL) (Maini et al., 2024; Shi et al., 2024), iv). Negative Preference Optimization on $D_{f}$ (NPO) (Zhang et al., 2024) and v). Representation Misdirection for Unlearning (RMU) (Li et al., 2024). These unlearning methods are detailed in Appendix A.2.
@@
Recent audits reveal that LLMs subjected to “unlearning” procedures may still leak private information. Membership-inference accuracy often rebounds on “forgotten” samples (Duan et al., 2024), and prompt-injection (Shumailov et al., 2024) can smuggle hidden instructions past guardrails, reviving passage-level content the model was meant to forget. Even fine-tuning on a handful of public texts can quickly restore excised knowledge (Hu et al., 2024). To address this issue, Xu et al. (2025b) integrates masking, distillation, and fact regularization techniques to resist inference; min–max optimization objectives are also employed (Fan et al., 2025); and noise-based data augmentation methods reframe forgettable tokens as backdoor triggers, enabling their systematic neutralization (Huu-Tien et al., 2025). Privacy-oriented work leverages token-specific training to suppress memorized tokens, halving membership-inference success rates (Tran et al., 2025), and fairness-centric studies show that reducing variance gaps can mitigate adversarial recovery of $D_{f}$ (Tran & Woo, 2025).
@@ 3.1 LLM Unlearning
We begin by formulating LLM unlearning. Let $D_{t}$ denote the original training dataset, from which a pretrained original model $f_{\text{O}}(\cdot;\theta)$ is obtained, where $\theta$ is the parameter of $f_{\text{O}}$. The goal of an unlearning algorithm $\mathcal{MU}$ is to remove the influence of a designated forget set $D_{f}$ from $f_{\text{O}}$ (Liu et al., 2025; Geng et al., 2025; Liu et al., 2024b). To ensure that the unlearning process does not significantly degrade the model’s overall utility, a retain set $D_{r}$ is typically introduced (Ren et al., 2025; Ji et al., 2024). In practice, the retain set $D_{r}$ and forget set $D_{f}$ are disjoint, i.e., $D_{r}\cap D_{f}=\emptyset$. Based on $D_{f}$ and $D_{r}$, $\mathcal{MU}$ typically defines two loss terms: a forget loss that penalizes residual influence from $D_{f}$, and a retain loss that encourages preservation of performance on $D_{r}$. These objectives capture the dual goals of forgetting and retention, and can be expressed as the following regularized optimization problem (Pan et al., 2025):
@@ 3.2 Relearning Attacks in LLM Unlearning
The robustness issue of LLMs unlearning is primarily reflected in the vulnerability of current methods to relearning attacks (Fan et al., 2025). These attacks aim to rapidly recover deleted knowledge by performing lightweight fine-tuning on the unlearned model $f_{\text{U}}$ using only a small number of samples from the forget set $D_{f}$. The attacker’s objective is as follows:
@@
where $\delta$ denotes the adversarial update to the $f_{\text{U}}$’s parameters, $D_{sub_{f}}\subset D_{f}$ is a small subset of $D_{f}$ in the attack, $\ell_{\text{relearn}}$ is the relearning objective, which is often defined to counteract the unlearning process, such as the general fine-tuning loss or the negative of the part of forget loss on $D_{f}$.
@@ 4.1 Motivation
To illustrate the robustness vulnerability in current LLM unlearning methods and establish the foundation for our approach, we conduct a systematic analysis of the relearning attacks. We first perform unlearning on the Zephyr-7B-beta model (Tunstall et al., 2023) using the standard Gradient Ascent (GA) method on the WMDP Bio dataset (Li et al., 2024). After unlearning, we simulate an adversary who fine-tunes the unlearned model for only two epochs on small subsets of the original forget set $D_{f}$ (40 samples). Performance is evaluated on the WMDP Bio QA test set, where lower accuracy indicates stronger unlearning. As shown in Figure 1(a), GA-based unlearning reduces accuracy from 64.45% (original model) to 27.10%, suggesting effective suppression of target knowledge. However, the relearning
@@
When $S_{\delta}(\theta)$ is large, even tiny parameter updates can drastically change the loss. In such sharp regions, adversaries need only a few fine-tuning steps on small subsets of $D_{f}$ to reverse unlearning, directly explaining the effectiveness of small-shot relearning attacks. We visualize the loss surface around unlearned parameters by scanning two orthogonal directions $\mathbf{r}_{1},\mathbf{r}_{2}$ and plotting $z=\ell(\theta+\alpha\mathbf{r}_{1}+\beta\mathbf{r}_{2})$. As shown in Figure 2, GA-based unlearning forms a sharp basin, while our method StableUN produces a flatter basin with lower neighborhood sensitivity, thereby resisting small-shot relearning. This analysis highlights the need to explicitly control $S_{\delta}(\theta)$ during unlearning:
@@ 4.2 Overview
To address these robustness limitations while explicitly balancing the dual objectives of forgetting and remembering, we propose a feedback‑guided unlearning framework. By injecting task‑specific guidance signals into the optimization loop of mainstream LLM unlearning algorithms, the overall procedure becomes more robust and stable. Specifically, we design two kinds of feedback as follows:
  • The first is the forgetting feedback. We introduce a robustness‑oriented feedback signal that exposes the model to a family of parameter perturbations. Motivated by the idea that relearning attacks essentially act as weight‑space disturbances, we require the unlearned model $f_{\text{U}}$’s performance on the forget dataset $D_{f}$ to remain invariant to small perturbations on the parameter space, thereby enhancing its resistance to relearning attacks.
@@
Inspired by meta-learning (Andrychowicz et al., 2016; Wang et al., 2020; 2021), we formulate our approach as a bi-level optimization problem: an inner-loop unlearning step produces a temporary model, while an outer-loop feedback step refines the final update direction. This formulation introduces two stages into the standard unlearning loop, namely the unlearning-tuning stage and the feedback stage. For the unlearning-tuning stage, we create a temporary model $f_{\text{tmp}}(\cdot;\theta^{\tau})$ by running one gradient update with the unlearning loss (e.g., GA, NPO, RMU), which provides the basic direction to forget. Our feedback signals build on top of it. For the feedback step, the temporary model is then probed by the two complementary feedback signals introduced above, generating loss terms $\mathcal{L}_{f}$ and $\mathcal{L}_{r}$, respectively. Depending on the feedback target, the two signals respectively reflect the model’s forgetting performance within its current neighborhood and utility, revealing how well it robustly forgets $D_{f}$ and retains knowledge from $\hat{D}_{r}$. After these two stages, the algorithm executes the final cumulative parameter update using the harmonized gradient direction.
Given that the forgetting objective $\mathcal{L}_{f}$ and remembering objective $\mathcal{L}_{r}$ can induce conflicting update directions, we further introduce a gradient harmonization strategy to resolve this dilemma. Specifically, a simple projection operation removes the component of one gradient that conflicts with the other, producing a unified update direction that simultaneously promotes thorough deletion of $D_{f}$ and faithful preservation of essential knowledge in $D_{r}$. In the following sections, we detail our complete framework of StableUN.
@@ 4.3 Forgetting Feedback
To achieve a more thorough forgetting of the forget dataset $D_{f}$, especially to enhance the model’s robustness against relearning attacks, we propose forgetting feedback by applying various perturbation techniques to the model parameters to simulate the relearning attacks, ensuring the unlearned model’s performance on $D_{f}$ remains stable. The detailed procedure is as follows: As shown in Figure 3(a), the first step of forgetting feedback is the unlearning-tuning step. Initially, we construct a temporary model $f_{\text{tmp}}(\cdot;\theta^{\tau})$ by performing one gradient update using a standard forgetting loss (e.g., GA, GA+GD, GA+KL, NPO, RMU) on the original model $f_{\text{O}}$’s parameters $\theta$:
@@
The first category is chosen because it deliberately pushes the model toward worst-case directions, thereby simulating adversarial attempts to “re-awaken” forgotten information. This group includes Sharpness-Aware Perturbation (SAP(Foret et al., 2020), which perturbs parameters along the normalized gradient direction to capture the steepest ascent of local sharpness. It explicitly targets the direction that maximally increases loss, which is a typical adversarial move; Gradient-Aligned Perturbation (GAP(Moosavi-Dezfooli et al., 2019), which scales the gradient to drive the model into high-curvature regions; and Historical Weight Smoothing (HWS(Izmailov et al., 2018), which averages the current weights with several past checkpoints to test whether smoothing can implicitly roll parameters back toward memorized regimes. The second category introduces stochastic disturbances, represented by Gaussian Parameter Noise (GPN(Cohen et al., 2019), which adds unbiased Gaussian noise to the parameters and serves as a non-adversarial baseline.
Together, these perturbations form a spectrum from random to adversarial, providing a comprehensive probe of how resilient the unlearned model remains to relearning attempts. Their mathematical details are provided in Appendix A.3. Then we get the forgetting feedback. In each iteration, we randomly select $T$ perturbation methods from the above candidates to generate corresponding perturbed parameters $\{\theta^{\tau^{\prime}_{i}}\}_{i=1}^{T}$. These perturbed models are then evaluated on the relevant dataset(s) according to the specific $\mathcal{MU}$ method being used. The final forgetting feedback loss is as follows:
@@ 4.4 Remembering Feedback
To prevent unintended loss of useful knowledge in $D_{r}$ and preserve downstream utility, we introduce remembering feedback as a balance term, which parallels the forgetting branch but focuses on retention. After the standard unlearning tuning step (Eq. 5), we obtain a temporary model $\theta^{\tau}$. We then conduct a feedback-evaluation step on a query set $Q$, randomly sampled from a small retained subset $\hat{D}_{r}\subset D_{r}$ or a public corpus, to monitor utility degradation. The cross-entropy over the $M$ queries is averaged to form the remembering feedback loss, which penalizes drops in performance and guides optimization toward preserving generalizable knowledge while erasing $D_{f}$:
@@
where $\lambda_{r}$ modulates the strength of the remembering feedback. The formulation enables our feedback-guided remembering mechanism to penalize any degradation of informative samples.
@@
At each training step, StableUN produces two distinct feedback signals: a robustness-oriented forgetting loss $\mathcal{L}_{fb}$ and a utility-preserving remembering loss $\mathcal{L}_{fb_{2}}$. Combining these signals with the base forgetting objective yields the overall optimization target:
@@
We then compute their inner product $\langle g_{f},g_{r}\rangle=g_{f}^{T}g_{r}$. If this value is negative, the two directions conflict, and we project $g_{f}$ onto the sub-space orthogonal to $g_{r}$; otherwise, we keep $g_{f}$ unchanged:
@@ 5.1 Experiment Setup
We conduct experiments on two LLM unlearning benchmarks: (1) The WMDP benchmark (Li et al., 2024), which evaluates unlearning capabilities in hazardous domains such as biosecurity (WMDP-bio), cybersecurity (WMDP-cyber), and chemical safety (WMDP-chem). We primarily focus on the first two; (2) The MUSE benchmark (Shi et al., 2024), which contains two unlearning scenarios: News and Books. The former seeks to unlearn knowledge related to BBC news articles, while the latter aims to unlearn text fragments from the Harry Potter book series. Following previous literature and to demonstrate the effectiveness of our method across different models, we employ Zephyr-7B-beta (Tunstall et al., 2023) as the initial model for WMDP-bio, Mistral-7B (Jiang et al., 2023) fine-tuned on cybersecurity datasets as the initial model for WMDP-cyber, LLaMA-2 7B (Touvron et al., 2023) fine-tuned on BBC news for MUSE News, and ICLM 7B (Fan et al., 2024) fine-tuned on Harry Potter books for MUSE Books.
@@
As illustrated in §2.1, we use 5 methods: GA, GA+GD, GA+KL, NPO, and RMU. More details are included in Appendix A.2. The performance of LLM unlearning is evaluated through $\mathcal{MU}$ Effectiveness and model Utility Retention. For WMDP, $\mathcal{MU}$ effectiveness is measured by accuracy on the WMDP test set, where lower accuracy indicates better unlearning performance. Utility retention is assessed through zero-shot accuracy on MMLU, while higher utility scores reflect better retention of general capabilities. For the MUSE dataset, following (Shi et al., 2024), we measure performance through Verbatim Memory (VerbMem) and Knowledge Memory (KnowMem) on $D_{f}$, where lower values indicate better unlearning effectiveness. VerbMem compares the average ROUGE-L F1 score (Klimt & Yang, 2004) between model-generated continuations and ground-truth continuations for unlearning samples, while KnowMem compares the average ROUGE score between model responses and ground-truth answers for question-answer pairs in the $D_{f}$. Utility is computed through KnowMem on the retained set, calculated as the average ROUGE score for question-answer pairs on the retained set. We further evaluate the robustness under two adversarial scenarios: 1) Relearning attacks (Hu et al., 2024): This constitutes our primary research focus. We randomly sample relearning data from the $D_{f}$, with results averaged over 5 independent random trials. 2) Jailbreaking attacks (Ma et al., 2024; Łucki et al., 2025; Thompson & Sklar, 2024): We employ an enhanced GCG algorithm (Łucki et al., 2025) to achieve unlearning knowledge extraction through the generation of adversarial prefixes.